OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). (sin x)(cos x) = 0

OpenStudy (anonymous):

Explain how to do it please??

OpenStudy (anonymous):

Do you have Multiple choice answers?

OpenStudy (anonymous):

pi/2, π 0, pi/2, π, 3pi/2 π, 3pi/2 0, 3pi/2

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

So. With this problem, you are already at the final step, and all you have to do is match them up on a unit circle. Can you open one please? (:

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

Okay let me start here. \[(\sin(x))(\cos(x))=0 \] from this, to solve for both Sin and Cos you make \[Sin(x)=0\] and the \[Cos(x)=0\]

OpenStudy (anonymous):

Do you get that?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Alright and after that you pull out your handy dandy Unit circle

OpenStudy (anonymous):

Yep i have one

OpenStudy (anonymous):

You know on the unit circle there is (x,y) correct? well on a unit circle the x=cos and the y=sin.

OpenStudy (anonymous):

i know that.

OpenStudy (anonymous):

So if you look at the equations we just had. \[Sin(x)=0\] and \[Cos(x)=0\] where on the unit circle does Sin=0 and where does Cos=0

OpenStudy (anonymous):

sin=0 at0 0,1 and 0,-1

OpenStudy (anonymous):

cos=0 at 1,0 and -1,0

OpenStudy (anonymous):

oh other way around lol

OpenStudy (anonymous):

so sin=0 at 1,0 and -1,0 and cos=0 at 0,1 and 0,-1

OpenStudy (anonymous):

Yup :P But what are the Radian measures of that? (i.e \[\frac{ 5\Pi }{ 6 }\]

OpenStudy (anonymous):

oh right right.... OHHH those must be the answers right *epiphany* lol 0, pi/2, pi, 3pi/2?

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

Yup :D which is you second answer. Good job ^^

OpenStudy (anonymous):

Thanks and thank you for your help :D

OpenStudy (anonymous):

No problem. If you ever need anymore help let me know(: